# All Optical Echo State Network Reservoir Computing

### v2

@@ Abstract
We propose an innovative design for an all-optical Echo State Network (ESN), an advanced type of reservoir computer known for its universal computational capabilities. Our design enables fully optical implementation of arbitrary ESNs, featuring flexibility in optical matrix multiplication and nonlinear activation. Leveraging the nonlinear characteristics of stimulated Brillouin scattering (SBS), the architecture efficiently realizes measurement-free nonlinear activation. The approach significantly reduces computational overhead and energy consumption compared to traditional software-based methods. Comprehensive simulations validate the system’s memory capacity, nonlinear processing strength, and polynomial algebra capabilities, showcasing performance comparable to software ESNs across key benchmark tasks. Our design establishes a feasible, scalable, and universally applicable framework for optical reservoir computing, suitable for diverse machine learning applications.
@@ I Introduction
Recurrent neural networks (RNNs) have inherent memory capabilities, and are trained on sequential or time series data to create a machine learning model that attempts to make accurate time series predictions [1]. However, their reliance on multiple layers of tunable weights leads to high computational and training costs. A potential way to overcome this cost is with a recent paradigm of RNNs with fixed internal weights, known as reservoir computers (RCs). Reservoir computing has been of recent interest, with applications to time-series prediction tasks, classification tasks, and much more [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Unlike RNNs, which require large amounts of time and resources, RCs offer a much less computationally expensive method for machine learning.
@@
Optical platforms for computation and machine learning can, in principle, achieve more efficient computation in terms of energy and time [13, 14]. One important challenge in all-optical approaches is how to realize the nonlinear activation. For this, various nonlinear optical processes have been studied including the nonlinear behavior of SBS [15]. It is important to note that, while the implementations of all-optical reservoir computers have been reported [16], the realization of an all-optical RC with universal computation capability (i.e., universality) has not yet been demonstrated to our best knowledge.
@@
Universality refers to a computing machine’s ability to perform any computational task with any desired accuracy. In a supervised learning context, it specifically denotes the capacity to approximate any input-output relationship with arbitrary precision. Consequently, having a mathematically rigorous proof of universality is very desirable, as it ensures that the computing scheme can robustly handle a wide range of tasks. Arbitrary RCs are not guaranteed to obey the conditions necessary to be capable of universal computations. We present, for the first time to our best knowledge, a new approach representing a novel all-optical implementation of Echo State Networks (ESNs), which have recently been shown to form a universal approximating class of RCs [17].
@@
ESNs are a family of RCs that are formed from iteratively applying a nonlinear transformation after a fixed linear transformation applied at each input time step [18]. ESNs are similar to RNNs except that the linear weights given to the state of the system at each time step are fixed for the entire run. The family of ESNs as a whole is universal [17] which means that for a given target sequence and input, we can find a specific ESN configuration that can approximate the target function to any desired accuracy.
@@
Here, we propose an innovative way to realize an all-optical implementation of any desired ESN configuration. The advantages provided by an all-optical reservoir architecture, as stated earlier, are fast computation and better energy efficiency than large software alternatives. For this reason, there has been growing interest in recent years in optical implementations of the matrix multiplication stage of machine learning architectures [19, 20, 21]. Here, we propose an in-fiber method for optical matrix-vector multiplication, which is followed by an in-fiber solution for optical nonlinear activation inspired by Slinkov et. al [22] using SBS. This unique combination of measurement-free matrix multiplication and nonlinear activation allows for a feasible design of the optical ESN.
@@
The paper is organized as follows: Section II covers the basic concepts of ESNs and the theory behind the SBS linear amplifiers and activation function, Section III addresses our proposed design for the all-optical ESN, and Section IV compiles the performance of the ESN on various tasks that evaluate its universality, and compares it to a software ESN of the same size. The tunability of the nonlinearity of SBS interaction is also analyzed in this section. Finally, Section V contains the concluding remarks.
@@ II.1 Echo State Network
A reservoir computer is required to satisfy the uniform convergence property [17]. This property asserts that the reservoir state converges to a unique sequence after many time steps, which means that the initial state of the reservoir from the distant past has no effect on the current state. This property can be extended by including the continuity of $g(x_{k},u_{k})$ and result in the fading memory property [23]. The fading memory property implies that the dependence of the state vector of the reservoir $x_{k}$ on the input $u_{k_{0}}$ at a time step $k_{0}$ must diminish as $k-k_{0}$ tends to infinity. This allows for a universal, repeatable, and deterministic reservoir while still maintaining some memory.
An ESN [18] is a type of reservoir computer with a form given by
@@
where $A$ is a random but fixed matrix and $B$ is a random, fixed vector. The function $f(z)$ is a nonlinear activation function that acts on the input $z=Ax_{k}+Bu_{k}$. For $N$ nodes of the reservoir, the matrix $A$ has dimensions $N$ x $N$ and the vector $B$ has dimensions $N$ x $1$. Uniform convergence of an ESN is guaranteed by the condition that $A^{T}A<a^{2}I_{n}$ where $a$ is a number determined by the nonlinear function $f(z)$ [24]. Hence, the fading memory property is also guaranteed for a stable ESN, which leads to the universal approximation capability of this class of reservoir computers.
@@ II.2 Nonlinear Activation Mechanism
SBS is a third-order nonlinear scattering process. The optical pump wave interacts with the acoustic wave of the medium, which is stimulated by the presence of the pump and probe waves. This interaction amplifies the probe wave by depleting the pump. Writing the SBS interaction in terms of coupled powers of the pump and probe optical waves [25], we have
@@
We have defined $P_{\mathrm{s}}(L)=P_{\mathrm{out}}$. We also define $P_{\mathrm{s}}(0)=P_{\mathrm{in}}$. Here, $L$ is the length of SBS interaction. Equation (7) does not have an analytical solution but is solved approximately using numerical methods. Within the large pump approximation, which means that the pump power is much larger than the probe, the solution is linear and controlled by the pump power. When the pump power is made dependent on the input power the solution is nonlinear as shown by Slinkov et. al [22]. By amplifying and splitting the input with an appropriate ratio, we are able to optimize the nonlinear activation. We use these two behaviors to encode matrix multiplications and nonlinear activation in the optical ESN implementation.
The behavior of the transcendental Equation (7) has been simulated numerically, and is shown in Figure 1. Particularly, the nonlinear behavior of SBS looks very similar to a sigmoid function, which is often used as an activation function for machine learning applications. The sigmoid function is plotted for comparison in Figure 1(c). The linear behavior is tunable by varying the pump power and the nonlinear regime shows different behavior for different values of the $gL$ product given in the exponent of Equation (7), and also for different values of the pump splitting ratio. This allows for an optimizable activation function, which we expand on later in the paper.
@@ III Optical ESN Model
The optical ESN model is shown in Figure 2. It is designed in an all-fiber platform to make use of highly nonlinear fiber-based SBS for the all-optical activation function. According to the governing Equation (4), the design goal is to implement a matrix-vector multiplication followed by nonlinear activation.
Figure 2: All Optical ESN Design
The architecture aims to physically realize the ESN update Equation (1) at each time step by sequentially performing matrix-vector multiplications followed by nonlinear activation Equation (4). As shown in Figure 2, the input $u_{k}$ is encoded by a variable attenuator (VOA) at each time step $k$. The required $B^{\prime}$ vector elements are encoded using Brillouin amplifiers with appropriately tuned pump powers according to Equation (7) operating in the large pump regime, which gives a linear input-output relation as shown in Figure 1. The vector $B^{\prime}$ is chosen such that $AB^{\prime}=B$ to satisfy the ESN design in Equation (4). The values of the $A$ matrix are realized with a combination of a VOA and a Brillouin amplifier. This allows access to arbitrary (positive or negative) matrix elements, as discussed further below.
The input vector is combined with the optical reservoir using fiber beam combiners. These are coherent optical beam combiners, which introduces the possibility of reduced performance due to phase fluctuations. Fiber systems are known to have phase fluctuations due to vibrations, which necessitates the use of a phase lock loop to ensure coherent operation of the beam combiners. Electronic servo systems, such as the system proposed by Liang et al. [26] can be used to mitigate this issue.
The matrix multiplication is performed in a series of sub-steps that each time step $k$ is divided into. Each sub-step is denoted by the index $m$, and represents multiplication of the elements of the $x_{k}$ vector with each successive row of the matrix $A$. These multiplications are performed by appropriately tuning the array of VOAs for the inputs and the pump powers in the array of Brillouin amplifiers. As these amplifiers must be modulated every sub-step, the pump powers are also controlled by VOAs. We allow for negative elements in the ESN design matrices by representing the values of the state vector in dBm units.
While the Brillouin amplifiers implement only positive multiplication values (as the operation takes place in units of W), the accompanying VOAs can be tuned in each sub-step to achieve the desired negative value at the output in dBm units. This allows us to access negative matrix elements by using the appropriate range of optical powers below 0 dBm. Thus, we realized the ESN equations (1) - (4) all in dBm scale. This method is possible as the output is continuously monitored by the final detection step and the Brillouin amplifiers must be updated multiple times per input time step in order to time-multiplex the values of the $A$ matrix. While the elements of the $A$ and $B$ matrices are fixed constants, the amplifier gains and VOA losses are adjusted in each instance to convert the equations from dBm units into physically realizable gain and loss values.
As an example, if we have a specific input to output relation for this matrix multiplication given by $s_{i}=A_{ij}x_{j}$, where $x_{j}$ is 1-1 dBm and we wish to multiply $A_{ij}=2$, we would set our Brillouin gain to no gain, and set the VOA to attenuate the signal in order to achieve an output $s_{i}$ of 2-2 dBm. If the input $x_{i}$ is +1+1 dBm and we wish to multiply still the same value $A_{ij}=2$, we would set the VOA attenuation to no attenuation, and set the Brillouin gain to amplify the signal to make $s_{i}=+2$ dBm. The gain and attenuation values are calculated as Gain or Loss $10\mathrm{log}_{10}(\frac{P_{\mathrm{out}}}{P_{\mathrm{in}}})$. We can control the Brillouin gain by appropriately tuning the pump power, as explained above. In this way we are able to time multiplex the construction of the $A$ matrix by continuously monitoring the input and varying the VOA attenuation and Brillouin gain accordingly.
At first glance, this might seem like a limitation. However, the conversion from matrix elements in dBm scale to actual hardware parameters is straightforward and does not significantly increase implementation difficulty. We have accounted for this in simulations as well, as is discussed further in Section IV. This also requires the SBS process to respond quickly enough to modulations of the pump. The realistic gain modulation speed of fiber Brillouin amplifiers by modulating the pump power is typically limited to tens of MHz (e.g., up to around 50–100 MHz in silica fibers). This limitation arises from the intrinsic response time of the SBS process, which is governed by the phonon lifetime (approximately 10 ns) or equivalently the Brillouin gain linewidth (50–100 MHz). For example, recent advances in optical imaging [27] have made use of the quick response time of Brillouin amplifiers by modulating at over 1 MHz. Modulating the pump amplitude faster than the limit set by the Brillouin gain linewidth would not allow the acoustic wave (phonons) to fully respond, resulting in reduced or ineffective gain modulation. This tradeoff is discussed further below.
The output from this matrix multiplication on the current sub-step is passed through a compensatory Brillouin amplifier to account for any losses during previous and successive fiber splitting and combining steps, and to pre-amplify the signal before the activation step. It is then passed to the nonlinear activation. The input to the activation function Brillouin amplifier must be used in a self-pumping scheme [22]. The previous amplification step allows for it to have sufficient power to pump the nonlinear interaction, giving a nonlinear function at the output as shown in Figure 1. Here, we also choose the ratio between the pump and input beams (the pump splitting ratio) appropriately to optimize the nonlinear shape of the activation function, which is expanded on in the next section.
The output is the $m$’th element of the state of the reservoir at the next time step, $k+1$. It is switched into the appropriate output branch by a $1:N$ switch and the output is given an appropriate delay and split before both being sent to a detector and back into the reservoir for the next time step. This design utilizes only a single nonlinear activation element, allowing for high scalability. After the necessary delays to line up all of the output vectors, and at the end of all of the sub-steps, the state vector has been successfully updated and fed back into the reservoir for the next time step. After collecting all of the state vectors in the detection steps, we use post-training according to Equation (2) to achieve the required estimated output in response to the time-series inputs $u_{k}$.
@@
This design allows for the choice of arbitrary matrix elements for the matrix $A$ and vector $B$, including negative elements in dBm units through the use of the appropriate attenuation before each Brillouin amplifier. If chosen according to the convergence properties of an ESN, this design can be used to implement an arbitrary ESN, which is a universal family of reservoir computers [17]. This means that given a desired accuracy for a task (such as prediction), this method can be used to realize an optical hardware that achieves that accuracy. The proposed design also allows for scalability, as the number of nodes is increased simply by increasing the number of parallel fibers running in the optical fiber reservoir. Although, scalability is naturally limited by available resources and energy.
An important consideration is the energy consumption of the system. While an arbitrary ESN can be implemented with this architecture, the power flowing through the system can approach unfeasible amounts. In order to limit this, the scaling criteria on the amplifiers and nonlinear activation is made stringent to ensure that the maximum value of $x_{k}$ in the system is $31$ dBm. The maximum pump power for the nonlinear activation required in this scenario is $3.2$ W for a pump splitting ratio of 1000. The pump splitting ratio is discussed further in Section IV. While this results in a slight degradation in performance, we must accept this tradeoff to keep the power budget low.
@@
The delay line solution adds a different kind of hardware complexity despite reducing the required number of Brillouin amplifiers. While this is a drawback of using a fiber-based approach, we consider this fiber-based architecture to be a preliminary step towards an on-chip all optical architecture. Choudhary et al. [28], and Yu et al. [29] have shown >50>50 dB of Brillouin amplifier gain in photonic chip and >8>8 dB of gain in thin-film lithium niobate solutions respectively, showing promise that a high-gain and tunable Brillouin amplifier for the construction of a similar on-chip optical ESN may be feasible in the near future. In addition, the on-chip photonic-circuit implementation will allow multiple nonlinear activating Brillouin amplifiers in parallel, obviating the need for delay lines. This future photonic-circuit implementation will ensure the scalability of the machine size.
On-chip implementations will also simplify the phase locking requirement. These systems have much lower phase fluctuations than fiber systems due to the solid on-chip structure, and approaches to combine coherent beams in Silicon On Insulator (SOI) fabrications, such as the method demonstrated by Jiao et al. [30], show promise for this ESN technology to be brought to an on-chip architecture.

### v3

# Optical Echo State Network Reservoir Computing

@@ Abstract
We propose an innovative design for an optical Echo State Network (ESN), an advanced type of reservoir computer known for its universal computational capabilities. Our design enables an optical implementation of arbitrary ESNs, featuring flexibility in optical matrix multiplication and nonlinear activation. Leveraging the nonlinear characteristics of stimulated Brillouin scattering (SBS), the architecture efficiently realizes measurement-free nonlinear activation. The approach significantly reduces computational overhead and energy consumption compared to traditional software-based methods. Comprehensive simulations validate the system’s memory capacity, nonlinear processing strength, and polynomial algebra capabilities, showcasing performance comparable to software ESNs across key benchmark tasks. Our design establishes a feasible, scalable, and universally applicable framework for optical reservoir computing, suitable for diverse machine learning applications.
@@ I Introduction
Recurrent neural networks (RNNs) have inherent memory capabilities, and are trained on sequential or time series data to create a machine learning model that attempts to make accurate time series predictions [hochreiter1997long]. However, their reliance on multiple layers of tunable weights leads to high computational and training costs. A potential way to overcome this cost is with a recent paradigm of RNNs with fixed internal weights, known as reservoir computers (RCs). Reservoir computing has been of recent interest, with applications to time-series prediction tasks, classification tasks, and much more [yan2024emerging, gauthier2021next, pathak2018model, vlachas2020backpropagation, zimmermann2018observing, canaday2018rapid, wang2024ultrafast, ehlers2025stochastic, ehlers2025improving, zhu2024practical, zhu2024minimalistic]. Unlike RNNs, which require large amounts of time and resources, RCs offer a much less computationally expensive method for machine learning.
@@
Optical platforms for computation and machine learning can, in principle, achieve more efficient computation in terms of energy and time [shastri2021photonics, wang2022optical]. One important challenge in optical approaches is how to realize the nonlinear activation. For this, various nonlinear optical processes have been studied including the nonlinear behavior of SBS [phang2023photonic]. It is important to note that, while the implementations of optical reservoir computers have been reported [duport2012all], the realization of an optical RC with universal computation capability (i.e., universality) has not yet been demonstrated to our best knowledge.
@@
Universality refers to a computing machine’s ability to perform any computational task with any desired accuracy. In a supervised learning context, it specifically denotes the capacity to approximate any input-output relationship with arbitrary precision. Consequently, having a mathematically rigorous proof of universality is very desirable, as it ensures that the computing scheme can robustly handle a wide range of tasks. Arbitrary RCs are not guaranteed to obey the conditions necessary to be capable of universal computations. We present, for the first time to our best knowledge, a new approach representing a novel optical implementation of Echo State Networks (ESNs), which have recently been shown to form a universal approximating class of RCs [grigoryeva2018echo].
@@
ESNs are a family of RCs that are formed from iteratively applying a nonlinear transformation after a fixed linear transformation applied at each input time step [jaeger2001echo]. ESNs are similar to RNNs except that the linear weights given to the state of the system at each time step are fixed for the entire run. The family of ESNs as a whole is universal [grigoryeva2018echo] which means that for a given target sequence and input, we can find a specific ESN configuration that can approximate the target function to any desired accuracy.
@@
Here, we propose an innovative way to realize an optical implementation of any desired ESN configuration. The advantages provided by an optical reservoir architecture, as stated earlier, are fast computation and better energy efficiency than large software alternatives. For this reason, there has been growing interest in recent years in optical implementations of the matrix multiplication stage of machine learning architectures [zhou2022photonic, xu202111, yang2013chip]. Here, we propose an in-fiber method for optical matrix-vector multiplication, which is followed by an in-fiber solution for optical nonlinear activation inspired by Slinkov et. al [slinkov2024all] using SBS. This unique combination of optical matrix multiplication and nonlinear activation allows for a feasible design of the optical ESN.
@@
The paper is organized as follows: Section II covers the basic concepts of ESNs and the theory behind the SBS linear amplifiers and activation function, Section III addresses our proposed design for the optical ESN, and Section IV compiles the performance of the ESN on various tasks that evaluate its universality, and compares it to a software ESN of the same size. The tunability of the nonlinearity of SBS interaction is also analyzed in this section. Finally, Section V contains the concluding remarks.
@@ II.1 Echo State Network
A reservoir computer is required to satisfy the uniform convergence property [grigoryeva2018echo]. This property asserts that the reservoir state converges to a unique sequence after many time steps, which means that the initial state of the reservoir from the distant past has no effect on the current state. This property can be extended by including the continuity of $g(x_{k},u_{k})$ and result in the fading memory property [boyd1985fading]. The fading memory property implies that the dependence of the state vector of the reservoir $x_{k}$ on the input $u_{k_{0}}$ at a time step $k_{0}$ must diminish as $k-k_{0}$ tends to infinity. This allows for a universal, repeatable, and deterministic reservoir while still maintaining some memory.
An ESN [jaeger2001echo] is a type of reservoir computer with a form given by
@@
where $A$ is a random but fixed matrix and $B$ is a random, fixed vector. The function $f(z)$ is a nonlinear activation function that acts on the input $z=Ax_{k}+Bu_{k}$. For $N$ nodes of the reservoir, the matrix $A$ has dimensions $N$ x $N$ and the vector $B$ has dimensions $N$ x $1$. Uniform convergence of an ESN is guaranteed by the condition that $A^{T}A<a^{2}I_{n}$ where $a$ is a number determined by the nonlinear function $f(z)$ [ehlers2023improving]. Hence, the fading memory property is also guaranteed for a stable ESN, which leads to the universal approximation capability of this class of reservoir computers.
@@ II.2 Nonlinear Activation Mechanism
SBS is a third-order nonlinear scattering process. The optical pump wave interacts with the acoustic wave of the medium, which is stimulated by the presence of the pump and probe waves. This interaction amplifies the probe wave by depleting the pump. Writing the SBS interaction in terms of coupled powers of the pump and probe optical waves [boyd2008nonlinear], we have
@@
We have defined $P_{\mathrm{s}}(L)=P_{\mathrm{out}}$. We also define $P_{\mathrm{s}}(0)=P_{\mathrm{in}}$. Here, $L$ is the length of SBS interaction. Equation (7) does not have an analytical solution but is solved approximately using numerical methods. Within the large pump approximation, which means that the pump power is much larger than the probe, the solution is linear and controlled by the pump power. When the pump power is made dependent on the input power the solution is nonlinear as shown by Slinkov et. al [slinkov2024all]. By amplifying and splitting the input with an appropriate ratio, we are able to optimize the nonlinear activation. We define this parameter as the “pump splitting ratio” $m$, given by
$m=\frac{P_{\mathrm{p}}}{P_{\mathrm{in}}}.$ (8)
We use these two behaviors to encode matrix multiplications and nonlinear activation in the optical ESN implementation. The behavior of the transcendental Equation (7) has been simulated numerically, and is shown in Figure 1. Particularly, the nonlinear behavior of SBS looks very similar to a sigmoid function, which is often used as an activation function for machine learning applications. The sigmoid function is plotted for comparison in Figure 1(c). The linear behavior is tunable by varying the pump power and the nonlinear regime shows different behavior for different values of the $gL$ product given in the exponent of Equation (7), and also for different values of the pump splitting ratio $m$, given by Equation (8). This allows for an optimizable activation function, which we expand on later in the paper.
@@ III Optical ESN Model
The optical ESN model is shown in Figure 2. It is designed in an all-fiber platform to make use of highly nonlinear fiber-based SBS for the optical activation function. In order to design the ESN update Equation 4, the required operations are addition and multiplication, followed by nonlinear activation. The nonlinear step is achieved by the self-pumped Brillouin amplification scheme as discussed above.
Figure 2: Optical ESN Design
We employ dBm units instead of linear units. In our algebra, all the power is converted to dBm units through
$P(\mathrm{dBm})=10\log_{10}(P(\mathrm{mW})/1\mathrm{mW}).$ (9)
For example, if we want to perform the following algebra in our unit:
$s_{\mathrm{out}}=x+u^{\prime},$ (10)
where $x$ is an arbitrary dimension vector and $u^{\prime}$ is a vector with the same dimension as $x$, we accomplish the above algebra in dBm units through the following physical realization:
$s_{\mathrm{out}}(\mathrm{mW})=x(\mathrm{mW})\cdot G,\quad G=10^{0.1u^{\prime}},$ (11)
where the operator $\cdot$ is the element-wise multiplication and $G$ is realized through the combination of Brillouin amplifiers and the Variable Optical Attenuators (VOAs) with the gain and the loss prescribed through the vector $u$. For the case of Brillouin amplifiers, $G$ is the tunable Brillouin gain, controlled by the pump power. In the large pump approximation applied to the transcendental Equation 7, it takes the form
$G=\mathrm{exp}(gLP_{\mathrm{pump}}).$ (12)
$G$ can also be set to negative values with the help of a VOA. Hence, we can adjust $P_{\mathrm{pump}}$ value to accomplish $G=10^{0.1u^{\prime}}$ (specifically, $P_{\mathrm{pump}}(u^{\prime})=(1/gL)(0.1\ln 10)u^{\prime})$.
The multiplication step uses somewhat different physical mechanism, with the help of an in-line optical power monitor. As an example, if we have a specific input to output relation for this matrix multiplication given by $s_{i}=A_{ij}x_{j}$, where $x_{j}$ is 1-1 dBm and we wish to multiply $A_{ij}=2$, we would set our Brillouin gain to zero gain, and set the VOA to attenuate the signal in order to achieve an output $s_{i}$ of 2-2 dBm. If the input $x_{i}$ is +1+1 dBm and we wish to multiply still the same value $A_{ij}=2$, we would set the VOA attenuation to zero attenuation, and set the Brillouin gain to amplify the signal to make $s_{i}=+2$ dBm. Hence, for given $A_{ij}$ parameter, implementing the appropriate combination of the required Brillouin gain and the amount of VOA attenuation accomplishes the multiplication operation in dBm units.
The full design is shown in Figure 2. The reservoir state vector $x_{k}$ (in dBm units) is implemented by the optical signal running through the optical fibers in parallel. The input $u_{k}$ is multiplied by the $B^{\prime}$ vector such that $AB^{\prime}=B$. Using our physical realization of the dBm unit addition scheme explained above, the vector $B^{\prime}u_{k}$ is added to the state vector by appropriately tuning the gain (or loss) of the Brillouin amplifiers (or VOAs). Please note that the combination of Brillouin amplifiers and VOAs implements both the addition and the subtraction of optical powers in dB units. This is followed by the matrix multiplication step:
$s_{\mathrm{out}}(\mathrm{dBm})=As_{\mathrm{in}}(\mathrm{dBm}),$ (13)
where $s_{\mathrm{in}}$ and $s_{\mathrm{out}}$ are vectors with dimension $N$ (the dimension of the state vector $x$), and $A$ is the matrix. This operation requires the following operation for each $i$’th entry of the vector:
$s_{\mathrm{out}}(i)(\mathrm{dBm})=\sum_{j=1}^{N}A_{ij}s_{\mathrm{in}}(j)(\mathrm{dBm}),$ (14)
which is the combination of the dBm unit addition and the dBm unit multiplication. We use the same device pairs of Brillouin amplifiers and VOAs while, first, performing the $A_{ij}s_{\mathrm{in}}(j)(\mathrm{dBm})$ multiplication operation in dB unit for all different $j$’s in parallel, and adding the $A_{ij}s_{\mathrm{in}}(j)$ in dBm units for all the $i$’s, one by one sequentially. Hence, we will require $N$ number of pairs of a Brillouin amplifier and a VOA for performing the $A_{ij}s_{\mathrm{in}}(j)(\mathrm{dBm})$ multiplications and one pair of a Brillouin amplifier and a VOA for adding $A_{ij}s_{\mathrm{in}}(j)$ in dBm units for all $j$’s sequentially.
We note that the accumulating addition steps are performed by switching the signal into a loop of fiber with the Brillouin amplifier and VOA pair, performing $\sum_{j}$ operation in individual time steps. This is followed by a single nonlinear activation element, to achieve each value of the output state vector. Thanks to this operation including switching into an output array of $N$ fibers with appropriate delays, the outputs from the matrix multiplication steps will be appropriately lined up into the next loop of the optical reservoir. The total number of the pairs of the Brillouin amplifier and a VOA is therefore $N+1$ as each $i$ entry requires 1 pair for dBm unit multiplication and one pair is used repeatedly for dBm unit addition.
After each row is accumulated, the resulting output is sent to the single nonlinear activation element. The pump is split off of the input with a ratio $1:m$ where $m$ is the pump splitting ratio defined by Equation (8). This implements the nonlinear activation function as shown in Figure 1. Any required compensatory amplification can also be achieved in the accumulation step before the nonlinear activation element. After the nonlinear activation, it is switched into the appropriate output fiber using a $1:N$ fiber switch. These outputs must be delayed to get an appropriately lined up parallel state vector. By tuning the pump splitting ratio $m$ and the $gL$ value of the self-pumped SBS function, the performance of the ESN can be optimized. This is explored in detail in Section IV.
To make this dBm scale matrix multiplication architecture more concrete, we provide a simple $2\times 2$ dimensional example. If the input vector $s_{\mathrm{in}}(=x_{k}+B^{\prime}u_{k})$ is given as $\left(\begin{smallmatrix}1\\[2.84526pt] 2\end{smallmatrix}\right)$ and the desired $A$ matrix is $\left(\begin{smallmatrix}1&-2\\[2.84526pt] 3&-1\end{smallmatrix}\right)$, the matrix multiplication $As_{\mathrm{in}}$ would be performed as follows: First, for the multiplication $\left(\begin{smallmatrix}1&-2\end{smallmatrix}\right)\left(\begin{smallmatrix}1\\[2.84526pt] 2\end{smallmatrix}\right)$, the in-line power monitors would be used to determine the input vector powers in dBm units. The first row $\left(\begin{smallmatrix}1&-2\end{smallmatrix}\right)$ is multiplied by the input vector $\left(\begin{smallmatrix}1\\[2.84526pt] 2\end{smallmatrix}\right)$ by allowing the 1 dBm power in the first node to pass unchanged realizing 1 dBm $(=1\times 1)$, and implementing 6-6 dBm of loss to the incoming 2 dBm power in the second node realizing 4-4 dBm $(=-2\times 2)$. The output values of (1,4)(1,-4) are detected by the second set of in-line power monitors. In the next step, taking place in a fiber loop accessed by a fiber switch, the addition of $1$ and 4-4 occurs through the dBm unit addition scheme that we explained above, by implementing 4-4 dB loss to the VOA on the incoming 1 dBm power. This process produces 3-3 dBm power. This first row output is allowed to pass by the fiber switch to the nonlinear activation step and switched into the first output row by the $1:N$ switch, where it encounters a delay loop. For the second row operation realizing $\left(\begin{smallmatrix}3&-1\end{smallmatrix}\right)\left(\begin{smallmatrix}1\\[2.84526pt] 2\end{smallmatrix}\right)$, the first node of 1 dBm power is amplified by 2 dB to realize $3$ dBm $(=3\times 1)$ and the second node of power 2 dBm is attenuated by 4-4 dB to realize 2-2 dBm ($(-1)\times 2)$. The output values of (3,2)(3,-2) are now realized in dBm scale. Then, the next step realizes the dBm unit addition of 3 dBm and 2-2 dBm. For this, a Brillouin amplifier gain of 3 dB amplifies the incoming 2-2 dBm power, which results in 1 dBm (=3+(2))(=3+(-2)) power. This output is similarly passed to the nonlinear activation step and switched to the output, where it emerges from the architecture synchronized with the first vector element computed earlier. Overall, we accomplished the matrix multiplication $\left(\begin{smallmatrix}1&-2\\[2.84526pt] 3&-1\end{smallmatrix}\right)\left(\begin{smallmatrix}1\\[2.84526pt] 2\end{smallmatrix}\right)=\left(\begin{smallmatrix}-3\\[2.84526pt] 1\end{smallmatrix}\right)$ in dBm units, and passed the output of the matrix multiplication through the optical nonlinear activation function.
@@
This design allows for the choice of arbitrary matrix elements for the matrix $A$ and vector $B$, including negative elements in dBm units through the use of the appropriate attenuation and Brillouin gain. If chosen according to the convergence properties of an ESN, this design can be used to implement an arbitrary ESN, which is a universal family of reservoir computers [grigoryeva2018echo]. This means that given a desired accuracy for a task (such as prediction), this method can be used to realize an optical hardware that achieves that accuracy. The proposed design also allows for high scalability, as the number of nodes is increased simply by increasing the number of parallel fibers running in the optical fiber reservoir, and the architecture uses only a single nonlinear activation element through the use of fiber switched delays. Although, scalability is naturally limited by available resources and energy.
@@
This also requires the SBS process to respond quickly enough to modulations of the pump. The realistic gain modulation speed of fiber Brillouin amplifiers by modulating the pump power is typically limited to tens of MHz (e.g., up to around 50–100 MHz in silica fibers). This limitation arises from the intrinsic response time of the SBS process, which is governed by the phonon lifetime (approximately 10 ns) or equivalently the Brillouin gain linewidth (50–100 MHz). For example, recent advances in optical imaging [shaashoua2024brillouin] have made use of the quick response time of Brillouin amplifiers by modulating at over 1 MHz. Modulating the pump amplitude faster than the limit set by the Brillouin gain linewidth would not allow the acoustic wave (phonons) to fully respond, resulting in reduced or ineffective gain modulation.
Another important consideration is the energy consumption of the system. While an arbitrary ESN can be implemented with this architecture, the power flowing through the system can approach unfeasible amounts. In order to limit this, the scaling criteria on the amplifiers and nonlinear activation is made stringent to ensure that the maximum value of $x_{k}$ in the system is $31$ dBm. The maximum pump power for the nonlinear activation required in this scenario is $3.2$ W for a pump splitting ratio of $m=1000$. The pump splitting ratio is discussed further in Section IV. While this results in a slight degradation in performance, we must accept this tradeoff to keep the power budget low.
Another important trade-off is between scalability and speed of operation involved. The current approach uses sequential time delays to simplify the hardware requirements. Alternatively, the architecture can be modified to fix the gain of the Brillouin amplifiers, with appropriate attenuation beforehand to always maintain convergence. This will result in a fixed reservoir implementation, but will no longer represent an ESN as the $A$ matrix will change at every time step due to a lack of correction for the input optical power. However, this comes at the cost of the guarantee of universal applicability for an ESN, therefore we have decided against this approach despite the relative hardware simplicity.
The sequential delay solution adds a different kind of hardware complexity despite reducing the required number of Brillouin amplifiers. While this is a drawback of using a fiber-based approach, we consider this fiber-based architecture to be a preliminary step towards an on-chip all optical architecture. Choudhary et al. [choudhary2016advanced], and Yu et al. [yu2025chip] have shown >50>50 dB of Brillouin amplifier gain in photonic chip and >8>8 dB of gain in thin-film lithium niobate solutions respectively, showing promise that a high-gain and tunable Brillouin amplifier for the construction of a similar on-chip optical ESN may be feasible in the near future. This future photonic-circuit implementation will ensure the scalability of the machine size.